OpenStudy (mandii):

1/sec x + 1/sec x = sec x

OpenStudy (anonymous):

\[\frac{2}{secx}=secx\]\[2=\sec^2x\]\[\frac{1}{\cos^2x}=2\]\[\cos^2x=\frac{1}{2}\]

OpenStudy (anonymous):

\[cosx= \pm \sqrt{\frac{1}{2}}\]

OpenStudy (kira_yamato):

\[\frac{2}{\sec x} = \sec x\] \[\sec^2x = 2\] \[\cos^2x = \frac{1}{2}\] \[\cos x = ±\frac{1}{sqrt {2}}\]

OpenStudy (anonymous):

\[x=\cos^{-1} \sqrt{\frac{1}{2}}\]or\[x=\cos^{-1}(-\sqrt{\frac{1}{2}})\]